# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ IV Commercial Applications of ZKML
Currently, some projects and companies have already begun using ZKML technology to protect the privacy and security of data for machine learning. This section introduces several companies and projects that have explored and implemented ZKML in practice.
EZKL [77] is a library and command line tool for doing inference for deep learning models using zk-SNARK. With EZKL, we first define a neural network model in Pytorch or TensorFlow, and then export the model as an ONNX file with some sample inputs in a JSON file, and finally apply EZKL to these files to generate a zk-SNARK circuit. With the latest round of performance improvements, EZKL can prove a MNIST-sized model in about 2 seconds and takes up 700MB of memory. So far, EZKL has been used as infrastructure in several hackathon projects, showing significant early adoption.
Coda [78] was the first cryptocurrency protocol with a clean blockchain. Current cryptocurrencies such as Bitcoin and Ether store hundreds of gigabytes of data, which will only grow larger over time. For Coda, however, no matter how much usage grows, the stored data always remains the same size, around 20 KB. By utilizing recursive zk-SNARK, Coda block producers can quickly share proofs of the correct blockchain state across the network and easily update the proofs as new transactions occur. This breakthrough application of zero-knowledge proofs enables Coda to provide scalability for thousands of transactions per second, millions of users, and years of transaction history without sacrificing security.
The DeFiChain[79] platform uses ZKML technology to validate the accuracy of AI models used for fraud detection and credit risk assessment without revealing any user data. DeFiChain uses zk-SNARKs to generate cryptographic proofs that their AI models produce accurate risk assessments. This process can be done without revealing the inner workings of the model, such as coefficients in logistic regression or decision trees in random forests, and assuring users that the model is working properly.
Modulus Labs[80] demonstrated ZKML use cases with RockyBot (an on-chain trading bot) and Leela vs. the World (a chess game). In particular, RockyBot is the world’s first fully on-chain AI trading bot, which brilliantly uses AI/ML for asset management. rockyBot uses the zk-SNARKs technique to demonstrate that the bot’s trading decisions agree with the predictions of trained Recurrent Neural Networks (RNN) and Long Short-Term Memory (LSTM) models. In the Leela vs. the World game, everyone can play against a verified on-chain instance of the Leela chess engine. The team has also written a paper [7] that benchmarks the speed and efficiency of various proving systems with different model sizes.
Giza [81] is a protocol that allows AI models to be deployed on-chain using a completely trustless approach. The technology stack used by Giza consists of the following four: 1) the ONNX format for representing machine learning models, 2) the Giza Transpiler for converting these models into the Cairo program format, 3) ONNX Cairo Runtime for executing the models in a verifiable and deterministic way, and 4) Giza Model smart contracts for deploying and executing models on the chain.
ZKaptcha [82] focuses on the problem of bots in web3 to provide CAPTCHA services for smart contracts. The current implementation of the project allows end users to generate proof that it is a human operation by completing a CAPTCHA. The CAPTCHA is validated by an on-chain validator and accessed through a smart contract. Currently, ZKaptcha relies primarily only on ZKP, but there are plans to implement ZKML in the future, similar to existing web2 CAPTCHA services. These services can analyze behaviors such as mouse movements to determine if a user is human, rather than a bot.
V Future Directions of ZKML
With the rapid development of ZKML technology, the future research and application directions present a broad prospect. Combining the current research status and technical challenges of ZKML, we conclude the following possible development directions:
  • Improve computational efficiency and scalability: Although existing ZKML frameworks have been able to handle simple machine learning models, computational efficiency and scalability are still key bottlenecks when dealing with large-scale deep neural networks, multimodal models, and generative models. Future research should focus on optimizing zero-knowledge proof protocols to support efficient proof generation and verification for large-scale computational tasks. For example, novel recursive proof systems, optimized arithmetic circuit designs, and the development of specific ZK protocols suitable for deep learning can be explored. This will pave the way for large-scale deployment of ZKML in real-world applications.
  • Optimize generalization and practicality: Multiple hurdles still need to be crossed to move ZKML technology from theoretical research to practical applications. Future research can be devoted to developing more practical ZKML tools and platforms, lowering the threshold of development and application, and making it widely used in various industries. For example, finance, healthcare, cryptocurrency, supply chain and other fields have a high demand for data privacy and security, and ZKML has a broad application prospect in these fields. Further research on how to integrate ZKML technology into existing machine learning frameworks to enhance its operability in industry is key to advancing this field.
  • Support for diverse and complex machine learning models: As machine learning models continue to become more complex, ZKML needs to extend the range of models it supports, especially emerging machine learning paradigms such as multimodal models, generative models, and self-supervised learning models. Future research could explore how to efficiently verify the correctness of these complex models in a zero-knowledge proof framework. For example, developing special proof techniques for generative models or zero-knowledge proof works for multimodal fusion. In addition, self-supervised and unsupervised learning models with complex loss functions and iterative and diverse optimization processes need to be better supported in ZKML.
  • Enhance privacy protection and security: Privacy protection is one of the core goals of ZKML, and future research needs to further enhance the privacy protection of models and data while ensuring the security of the models. Currently, most of the research focuses on privacy protection in the inference phase, and future research should also focus on privacy protection in the training phase as well as comprehensive protection of the entire machine learning pipeline. For example, research can be conducted on how to verify the correctness of the model training process while maintaining model privacy, or combine techniques such as homomorphic encryption and multi-party secure computation to enhance the security of models and data. In addition, the defense capability of the model against adversarial attacks should be strengthened by detecting adversarial samples through zero-knowledge proof techniques to ensure that the model maintains stable and trustworthy performance in the face of malicious attacks, thus enhancing the overall security of the model while protecting privacy.
VI Conclusions
This survey comprehensively investigates existing ZKML works. First, we introduce the backgrounds on machine learning and zero-knowledge proofs, followed by an exploration of zero-knowledge proof-based verifiable machine learning (ZKML). We place special emphasis on the unique advantages of zero-knowledge proof techniques over other cryptographic methods in terms of protecting privacy and verifying the computational correctness of machine learning. Subsequently, we categorize ZKML works according to the different stages of the machine learning process and the associated verification goals, while reviewing the current state of the art in research. We also discuss several commercial applications of ZKML. Finally, we propose future directions for the development of ZKML, with a particular focus on enhancing computational efficiency and scalability, optimizing performance, expanding application ranges, and improving privacy protection and security. These discussions and analyses of ZKML provide valuable insights and guidance for both academia and industry.

### v2

@@ IV Commercial Applications of ZKML
Beyond academic prototypes, a small but growing commercial ecosystem has emerged around ZKML. Current efforts fall into two broad categories: enabling toolchains that lower the barrier to deploying proof-backed ML, and application-facing platforms that use ZKP or ZKML to support trust-minimized decision-making, on-chain agents, or human-verification services.
EZKL is one of the most visible general-purpose toolchains for zkML inference [75]. In practice, models are exported from PyTorch or TensorFlow into ONNX, after which EZKL compiles them into zk-SNARK circuits. The framework also supports associated input formatting, witness generation, and proof creation. With continued engineering optimizations, it can prove MNIST-scale models within a few seconds while using modest memory by current zkML standards, making it important enabling infrastructure for prototypes and hackathon-style deployments.
Although not itself a ZKML platform, Coda illustrates the commercial maturation of recursive zk-SNARK infrastructure that later on-chain ML deployments can build upon [76]. Its constant-size blockchain proof shows how large computation histories can be compressed into succinct public proofs, a capability that is highly relevant to future verifiable ML services deployed in decentralized settings.
DeFiChain publicly highlights the use of ZKML-style ideas for AI-assisted fraud detection and credit-risk assessment [77]. The central motivation is to validate model outputs without exposing user data or proprietary model internals, thereby combining financial trust with privacy preservation.
Modulus Labs has been one of the most visible teams pushing on-chain ZKML demonstrations [78]. Its examples include RockyBot, an on-chain trading bot whose actions are tied to predictions from recurrent models, and Leela vs. the World, which exposes a verified on-chain chess engine. Modulus Labs has also published a benchmarking study comparing the speed and cost of different proving systems across model sizes [7], helping frame the engineering trade-offs of practical ZKML deployment.
Giza focuses on trustless on-chain model deployment [79]. Its stack combines ONNX for model representation, a transpiler from ONNX to Cairo, an ONNX-Cairo runtime for deterministic execution, and smart-contract components for deployment and interaction. This pipeline illustrates how model serving, proof-oriented execution, and blockchain integration can be combined into a unified deployment workflow.
ZKaptcha targets bot resistance in Web3 environments by providing CAPTCHA-style proof services for smart contracts [80]. Its current implementation relies mainly on ZKP rather than full ZKML, but it points to a natural adjacent application area: proof-backed behavioral classification that can distinguish humans from automated agents without exposing sensitive interaction data. Taken together, these examples suggest that early commercial adoption of ZKML is clustering around on-chain inference, verifiable agent logic, and trust-minimized user interaction.
V Hybrid Cryptographic Frameworks for Verifiable ML
In practice, ZKP is often most effective when combined with complementary privacy-enhancing technologies rather than used in isolation. Homomorphic encryption, differential privacy, federated learning, and secure multiparty computation each address a different part of the trust problem, but none of them alone provides full computational verifiability. Hybrid designs therefore aim to combine privacy, integrity, and scalability by assigning each primitive the role it handles best. This section highlights representative examples of such combinations.
V-A Hybridization of Homomorphic Encryption and ZKP
pvCNN [33] illustrates how homomorphic encryption and ZKP can play complementary roles in verifiable inference. Homomorphic encryption protects sensitive intermediate computation, while the proof system attests that the returned result is consistent with the prescribed CNN execution. Drynx [37] shows a broader hybrid pattern in distributed analytics by combining homomorphic encryption, differential privacy, and ZKP so that confidentiality, secure aggregation, and verifiability are handled by different layers of the system rather than by a single primitive.
V-B Hybridization of Differential Privacy and ZKP
Verifiable differential privacy [81] targets a different trust gap: even if a mechanism claims to satisfy differential privacy, users still need assurance that the stated noise distribution, sensitivity bound, and privacy budget were actually respected. The scheme therefore couples differential-privacy computation with ZKP-based verification of key steps such as noise generation and parameter compliance. Its main technical challenge is that common differential-privacy mechanisms are not naturally proof-friendly, so the work relies on circuit-oriented linearization techniques to express them within arithmetic constraints.
V-C Hybridization of Federated Learning and ZKP
VPFL [82] shows how federated learning can be strengthened by adding both privacy-preserving aggregation and verifiability. In this design, local updates are protected during aggregation with homomorphic encryption, while SNARK-style proofs attest that the server carried out the prescribed federated-averaging computation. The result is a hybrid architecture in which confidentiality of client updates and correctness of aggregation are handled separately but coherently. A complementary perspective is provided by Wang et al. [83], who organize zero-knowledge federated learning into a structured ZK-FL framework spanning different FL stages and tasks, and further propose Veri-CS-FL, in which clients prove the quality of local-model performance metrics before server-side selection. This line of work is useful because it highlights that ZKP can support not only aggregation integrity, but also upstream decision points such as trustworthy client selection.
V-D Hybridization of Secure Multiparty Computation and ZKP
CrypTen [84] is best viewed as a representative MPC substrate rather than a ZKP-integrated system. Its relevance here is that it clarifies the role MPC can play in hybrid verifiable ML: MPC can protect multi-party training or inference over private inputs, while ZKP can be layered on top to attest the correctness of selected computations, aggregates, or model-update rules. This separation of roles is appealing because it avoids forcing a single primitive to provide both high-throughput privacy and full public verifiability.
V-E Benefits of Hybrid Cryptographic Designs
The integration of zero-knowledge proofs with complementary cryptographic techniques offers several advantages over standalone approaches:
  • Computational efficiency: Hybrid approaches delegate bulk confidential computation to primitives such as HE or MPC while reserving ZKP for correctness checks. For example, pvCNN shows that separating confidential computation from proof generation can reduce end-to-end overhead.
  • Security completeness: Multi-technology fusion covers multiple threat dimensions at once. Drynx, for instance, combines differential privacy, homomorphic encryption, and ZKP to address individual privacy, computation confidentiality, and protocol integrity simultaneously.
  • Scalability: Carefully chosen combinations can support larger deployments than any single primitive alone. VPFL demonstrates this point by maintaining logarithmic verification overhead as the number of participants grows.
  • Reduced trust assumptions: Hybrid designs can remove trust assumptions left open by individual technologies. Examples include verifiable differential privacy, which reduces reliance on honest implementation of privacy parameters, and MPC-based ML pipelines whose critical outputs can be attested with ZKP.
VI Future Directions for ZKML
Although ZKML has advanced rapidly in recent years, the field is still far from mature. Based on the current research landscape and the remaining technical bottlenecks, we highlight five especially important directions for future work.
VI-A Improving Computational Efficiency and Scalability
Although recent systems have substantially improved performance, efficiency and scalability remain the most immediate barriers to practical deployment. This is especially true for large neural networks, multimodal models, and generative systems, whose circuit sizes, witness-generation costs, and memory footprints can still be prohibitive.
Future work should therefore continue to explore recursive proof composition, more efficient circuit layouts, hardware-aware witness generation, and proof systems tailored to layered ML computation. Progress along these lines will be essential if ZKML is to move from small and medium models toward genuinely large-scale deployment.
VI-B Improving Generality and Practicality
Generality and usability remain another major bottleneck. Many current systems are tightly coupled to specific proof backends, quantization rules, operator sets, or model families, and they often require substantial cryptographic expertise to use effectively.
Future progress will likely depend on better compilers, reusable operator libraries, domain-specific languages, and interfaces that fit naturally into mainstream ML toolchains. More standardized development stacks would lower the entry barrier for practitioners. They would also make ZKML easier to evaluate, compare, and deploy across application domains.
VI-C Support for Diverse and Complex Machine Learning Models
The scope of supported models also needs to expand. Much of the current ZKML literature still centers on supervised inference or relatively structured training settings, whereas modern ML increasingly involves multimodal pipelines, generative models, retrieval-augmented systems, and self-supervised objectives.
Future research should investigate which components of these systems are most meaningful and tractable to verify. For example, one promising direction is to prove deterministic sub-computations while handling stochastic steps through verifiable randomness or carefully scoped correctness claims.
For self-supervised and unsupervised learning in particular, new theory may be needed to define verification targets that are both cryptographically sound and practically useful.
VI-D Strengthening Privacy and Security
Privacy protection is one of the core motivations for ZKML, but existing work still concentrates heavily on inference-time guarantees. Broader end-to-end protection is needed across training, deployment, updating, and model governance.
Future work should therefore study how ZKP can be combined more effectively with MPC, HE, differential privacy, and related techniques to provide stronger privacy guarantees across the full ML lifecycle.
Another open problem is how to reason about robustness and security beyond basic correctness, including adversarial manipulation, model misuse, and malicious updates. Extending ZKML toward these stronger security goals remains challenging but increasingly important.
VI-E ZKP-Based Verification of Data Authenticity
Data authenticity is another underexplored direction. Most current ZKML work assumes that the relevant datasets are already legitimate and properly governed, yet in practice the provenance, authorization, and integrity of training or testing data can be just as important as the correctness of the model computation itself.
Future work in this area could include proofs of dataset provenance, authorization, preprocessing integrity, and version consistency, together with scalable mechanisms for large distributed datasets. Bringing such guarantees into ZKML pipelines would significantly strengthen the trustworthiness of privacy-preserving ML services.
VII Conclusion
This survey has reviewed the foundations, representative systems, implementation bottlenecks, commercial efforts, and hybrid cryptographic directions of zero-knowledge machine learning. Across the literature, a clear pattern emerges: ZKML is moving from proof-of-concept demonstrations toward more general, efficient, and practically deployable systems, but substantial challenges remain in scalability, model coverage, privacy across the full ML lifecycle, and data authenticity. We hope this survey provides a useful reference for researchers and practitioners working toward trustworthy, privacy-preserving, and verifiable machine learning services.